# Direct in-situ observations of wave-induced floe collisions in the deeper Marginal Ice Zone

### v1

@@ Abstract
Ocean waves propagating through the Marginal Ice Zone (MIZ) and the pack ice are strongly attenuated. This attenuation is critical for protecting sea ice from energetic wave events that could otherwise lead to sea ice break-up and dislocation over large areas. Despite the importance of waves-in-ice attenuation, the exact physical mechanisms involved, and their relative importance, are still uncertain. Here we present for the first time direct in situ measurements of floe-floe interactions under the influence of waves, including collisions between adjacent floes. The collision events we report are aligned with the incoming wave direction, and phase-locked to the wave signal, which indicates that the individual collisions we detect are wave-induced. The observations demonstrate that wave attenuation by wave-induced floe-floe collisions, which have been studied in idealized laboratory and field experiments in the outer MIZ, are indeed a likely source of wave energy dissipation deep inside the MIZ as well.
@@ 1 Introduction
The outermost zone of the polar ice sheets is called the Marginal Ice Zone (MIZ) and is often defined as the region where the sea ice concentration is between 15% and 80% (Aksenov et al., 2017; Strong and Rigor, 2013). In the MIZ, ocean waves are strongly attenuated, preventing break-up of the ice further in (Kohout and Meylan, 2008; Bennetts and Squire, 2012). The details of the mechanisms causing wave-in-ice attenuation are still unclear, and a number of processes have been suggested in the literature, including wave refraction and diffraction by individual floes (Bennetts et al., 2010), viscous damping at the ice-water interface (Zhao et al., 2015; Sutherland and Rabault, 2016; Sutherland et al., 2019), viscoelasticity (Zhao and Shen, 2018; Zhang and Zhao, 2021), and floe-floe interactions (Løken et al., 2022; Herman, 2018). Uncertainties related to waves-in-ice damping mechanisms and intensity are currently limiting the accuracy of coupled numerical weather prediction and climate models of the polar regions, and significant efforts are now going into improving their accuracy (Boutin et al., 2020). This uncertainty stems both from the large spread in different waves-in-ice attenuation parametrizations and the significant discrepancies with in-situ observations (Voermans et al., 2021).
@@
Collisions between ice floes have been conjectured to be an important type of floe-floe interaction, yet their existence has remained controversial in the academic community. In particular, Li and Lubbad (2018) observed in laboratory experiments that collisions can take place and induce significant wave energy dissipation. Smith and Thomson (2020) observed collisions between ice pancakes in the newly formed MIZ area. Rabault et al. (2019) used optical methods to observe collisions between ice chunks and the associated water pumping in a small scale laboratory experiment. This demonstrated that floe-floe collisions can cause wave-in-ice energy dissipation due to both the inelastic energy losses in the collisions themselves, as well as the increased water turbulence levels and effective eddy viscosity induced in the water immediately underneath the ice. Following these results, the study by Løken et al. (2022) found in idealized field experiments that floe-floe collisions can amount to up to 45 % of the energy dissipated when ice floes move close to each other, and confirmed that collisions, when occurring, are at the origin of complex water jets and eddies. Noyce et al. (2023) recently reported time series of floe-floe collisions in the outer Antarctic MIZ during a polar cyclone. Even though these observations were made during an extreme weather event and the sea ice concentration was moderate, and though ice floes were smaller than deeper in the MIZ, this still suggests that floe-floe interactions and collisions are important mechanisms in real world conditions. Floe-floe collisions are also believed to be strongly related to the floe size distribution and likely to play a role in the merging and breaking of ice floes (Shen and Ackley, 1991; Herman, 2018). Recently, Rabault et al. (2024) has presented observations that may indicate that sea ice convergence could modulate the intensity of floe-floe collisions, with large impact on wave in ice attenuation.
@@
Here, we provide the first direct evidence that floe-floe interactions and collisions regularly happen deep into the MIZ and should be considered likely also under conditions representative of the deeper MIZ. We find more than 60 acceleration residuals in the horizontal plane that lie outside the 99.9% probability region in two time series. This corresponds to more than five times the expected amount of extreme events, and is analogous to a $3\sigma$-event in the one dimensional setting. These collisions are aligned with the dominating wave direction, and phase locked with the wave signal, strongly suggesting that the individual collisions we detect are wave-induced.
@@ 2.1 Buoy deployment and raw data recovery
The OMB-v2018 performs wave measurements using a Vectornav VN100 Inertial Motion Unit (IMU), and uses GPS to determine its position. The wave motion data are sampled over 20-minute intervals every four hours from records of vertical acceleration sampled at 10 Hz by the VN100 (Rabault et al., 2017). The VN100 has built-in accelerometer, gyroscope, and magnetometer, and performs on-board Kalman filtering (Rabault et al., 2016, 2020). This filtering ensures that the full 3-dimensional acceleration experienced by the buoy in the North-East-Down (NED) frame of reference is captured in the time series. Due to the presence of large batteries and cables carrying electrical currents in close vicinity to the VN100, the magnetometer data come with significant uncertainty and orientation relative to north should only be considered as an indication.
@@
Figure 2: Trajectories of the buoys deployed 2020-08 over the Yermak plateau area, superimposed on synthetic aperture radar (SAR) images of the MIZ on 2020-08-29. The buoy that was recovered is the northernmost one, which is shown in red. The markers indicate the position of each buoy on the 2020-08-29 at 12 UTC. The figure is created using the Ocean Visual Laboratory tool (permalink: https://odl.bzh/vEFWnkH0). Sea ice is clearly visible in the SAR images, and the buoy recovered is deep in the MIZ.
@@ 2.2 Signal processing of recovered in-situ time series
where $f_{\text{max}}$ is the frequency at which the PSD reaches its maximum value.
@@ 2.3 Detection of Collisions
The measured accelerations contain three contributions: (i) the motion of the ice floe the buoy is resting on while it follows the waves, (ii) other physically meaningful signals, in particular, as we will demonstrate in the following, floe-floe collisions, (iii) sensor and signal processing (in particular Kalman filter) noise. The latter two, i.e. (ii) and (iii), correspond to the residual signal computed above.
The noise contribution (iii) arises from the noise in the micro-electromechanical systems (MEMS) sensors used in the IMU, and its propagation through the embedded Kalman filter run on-the-fly. According to the datasheet of the IMU, and in agreement with previous observations (Rabault et al., 2020), this generally follows a normal distribution. As a consequence, statistically significant deviations from a normal distribution in the residuals is a likely signature that physically meaningful signals, corresponding to actual signal (ii) rather than noise (iii), are happening. To detect this, we compute the 99% and 99.9% probability areas for the residuals under an assumption that the residuals are normally distributed. If the residuals deviate from these probability areas, the residuals are most likely not normally distributed and likely contain physically meaningful signal (ii).
@@ 2.4 Ocean and sea ice model data
Model data leveraging assimilation of satellite observations are used to estimate ocean and sea ice conditions in the area where the recovered buoy was active. In particular, we use the WAM-4km model data, run operationally at the Norwegian Meteorological Institute (MetNo), to estimate the open ocean wave conditions. WAM-4km is run using WAM Cycle 4.7. provided by the Helmholtz-Zentrum Geesthacht (Group, 1988; Günther et al., 1992). The version used at MetNo has several updates. The most relevant are: i) new physics based on Ardhuin et al. (2010), ii) correcting the growth of the waves in very high winds (Breivik et al., 2022). The corresponding model data are openly available [^2] https://thredds.met.no/thredds/catalog/fou-hi/mywavewam4archive/catalog.html.
Sea ice conditions estimates are gathered from the Nordic4 model, run operationally at MetNo and openly available [^3] https://thredds.met.no/thredds/catalog/fou-hi/nordic4km-zdepths1h/catalog.html. The Nordic4 model uses the simple Budgell sea-ice model (Budgell, 2005), with flux correction and nudging of the sea ice concentration towards observed ASMR2 sea ice concentration (Spreen et al., 2008). This provides access to robust estimates of both the sea ice concentration (SIC) and sea ice thickness (SIT).
@@ 3.1 Collisions detection: direction and phase locking with incoming waves
The extreme values in the residuals are likely to be caused by floe-floe collisions induced by the incoming waves. This is supported both by the general alignment of the residual acceleration extreme events in the horizontal plane with the dominating incoming wave direction, and the absence of extreme residuals in the vertical direction. In order to further cross-check this hypothesis, we look at the correlation between the instantaneous wave phase obtained by the Hilbert transform of the double integrated wave elevation $\eta(t)$, and the occurrence of extreme events in the acceleration residuals. If the extreme residual events are independent of the wave field, and hence not wave driven collisions, we would expect a uniform distribution of extreme residuals over the phase of the wave field. As visible in Fig. 7, the extreme events are clearly not uniformly distributed, and hence the extreme events follow the phase of the wave field.
Figure 7: Phase diagram of the extreme residual values detected the two FOIs, relative to the incoming wave phase. The phase of the wave field is found using a Hilbert transform. Then, each extreme residual event (meaning it is outside the 99.9% probability region) is plotted at its respective phases. The upper half of the circle is the wave crest, while the bottom half is the wave trough. Clear phase clustering is observed in both FOIs, indicating that the occurrence of the extreme events we detect is strongly correlated to specific portions of the incoming wave phase motion.
@@ 3.2 Waves and sea ice conditions when collisions happen
Figure 8: Sea ice concentration (SIC) encountered by waves propagating into the MIZ following the peak wave direction, according to the sea ice model. The blue dashed line shows the threshold corresponding to a SIC of 5%. The red dots indicate the buoy location at the time corresponding to each SIC profile. All SIC profiles are aligned to start right at the start of the MIZ (according to the 5% SIC threshold criterion) at the 0 km distance. The two FOIs are highlighted in a dark color. As visible, the buoy is in the deep MIZ during the time corresponding to both FOIs.
@@
Figure 9: Situation map showing the sea state for the two FOIs at the time when the collisions occurred. The sea ice concentration is shown in all four figures. The top row (figures (A) and (B)) shows the model swell partition significant wave height and direction at the time of F291 (left) and F312 (right), respectively, while the bottom row (figures (C) and (D)) similarly shows the model wind sea partition significant wave height and direction for the two FOIs. The position of the buoy at the time corresponding to each file is marked with a light blue star.
@@
Figure 10: Significant wave height (red) and peak period (blue) obtained from processing of the raw data files obtained from the SD card of the recovered buoy as a function of time. The green dashed lines indicate the times for which we detect collisions following the 99% confidence test and where we also observe clear signs of collisions in the raw acceleration. All time series with a $T_{p}>30$ seconds are marked with a lighter colouring to indicate files where there is little to no wave activity and the peak period saturates due to the low frequency noise inherent to acceleration measurements.
@@ 4 Discussion
While we have convincing evidence that our records contain wave-induced floe-floe collisions, it is difficult to be more specific about the detailed mechanics of these collisions. While the motion of the sea ice under the influence of waves is mostly happening in the vertical direction, some residual horizontal motion can also take place. Therefore, if neighboring ice floes are close enough, it is possible that this residual horizontal motion can lead to floe-floe collisions, as recently discussed in Rabault et al. (2024). While one could expect that adjacent floes with similar shapes move in (near) synchronization under the influence of monochromatic waves, real world conditions can lead to the existence of significant relative motion between floes. In particular, real world wave fields are not monochromatic and individual waves have a spread in both their individual amplitude and wavelength. This naturally leads to different forcing between consecutive waves, which can result in different motion between adjacent floes at any given time, and, hence, collisions. In addition, ice floe sizes typically follow a relatively broad floe size distribution (Toyota et al., 2006). This means that adjacent floes can have different responses to the same incoming waves, which, in turn, can also be a source of collisions. However, more work will be needed to investigate the detailed underlying mechanics, possibly by combining buoy observations with other more exhaustive techniques.
@@
The present study owes a lot to luck and serendipity. Both the fact that the instrument v-2018 contains a SD card with a full copy of the time series (which is due to the fact that the microcontrollers available back then were not able to perform the signal processing needed, so that an additional RaspberryPi had to be added for processing the time series), and the fact that the instrument recovered got stranded, was found, and that the SD card was not damaged, are a strike of luck. However, it would be easily possible, based on the experience obtained from the present dataset, to adapt existing open source instruments firmware, such as the OMB-v2021, to i) save raw time series on an embedded SD card in case the instrument is found again, ii) compute in-situ a number of collision statistics, such as the ones used here, and append these to the wave spectra that are being transmitted. Since a simple metrics of collisions, such as the number and relative strength of extreme acceleration events, and their correlation with the wave phase, would be only a few scalars, this would not add prohibitive transmission costs nor energy consumption to the design. Moreover, all the data needed to perform such an analysis are already available on-board. Adding such metrics to wave in ice buoys transmissions would allow to perform large scale, statistically representative analysis of the correlation between collision-like proxy metrics, changes in the sea ice and wave in ice conditions, and the wave in ice attenuation rate.

### v3

@@ Abstract
Ocean waves propagating through the Marginal Ice Zone (MIZ) and the pack ice are strongly attenuated. This attenuation is critical for protecting sea ice from energetic wave events that could otherwise lead to sea ice break-up and dislocation over large areas. Despite the importance of waves-in-ice attenuation, the exact physical mechanisms involved, and their relative importance, are still uncertain. Here we present direct in situ measurements of floe-floe interactions under the influence of waves, including collisions between adjacent floes. The collision events we report are aligned with the incoming wave direction, and phase-locked to the wave signal, which indicates that the individual collisions we detect are wave-induced. The observations indicate a possible correlation between sudden increases in wave activity and the frequency of floe-floe collisions.
@@ 1 Introduction
The outermost zone of the polar ice sheets is called the Marginal Ice Zone (MIZ) and is defined as the area along the ice pack that is affected by open ocean processes Wadhams et al. (1986); Barber et al. (2015). In the MIZ, ocean waves are strongly attenuated, preventing break-up of the ice further in (Kohout and Meylan, 2008; Bennetts and Squire, 2012). The details of the mechanisms causing wave-in-ice attenuation are still unclear, and a number of processes have been suggested in the literature, including wave refraction and diffraction by individual floes (Bennetts et al., 2010), viscous damping at the ice-water interface (Zhao et al., 2015; Sutherland and Rabault, 2016; Sutherland et al., 2019), viscoelasticity (Zhao and Shen, 2018; Zhang and Zhao, 2021), turbulence production and dissipation (Smith and Thomson, 2019), and floe-floe interactions (Løken et al., 2022; Herman, 2018). Uncertainties related to waves-in-ice damping mechanisms and intensity are currently limiting the accuracy of coupled numerical weather prediction and climate models of the polar regions, and significant efforts are now going into improving their accuracy (Boutin et al., 2020). This uncertainty stems both from the large spread in different waves-in-ice attenuation parametrizations and the significant discrepancies with in-situ observations (Voermans et al., 2021).
@@
Collisions between ice floes is believed to be an important type of floe-floe interaction, yet their relative importance as an attenuation mechanism, and the exact nature on how these collisions attenuate waves has been a topic of discussion in the academic community. These collisions were first observed in a series of field measurements in the Bering strait and off the coast of Greenland, see Martin and Becker (1987) and Martin and Becker (1988), but the phenomenon is seldom discussed in the literature. The literature on collisions was summarized in section 2.2 in Squire et al. (1995), while some mechanisms for wave attenuation was further summarized in the review articles Squire (2007) and Squire (2020), although the latter two of these reviews does not discuss floe-floe collisions in detail. In particular, Li and Lubbad (2018) observed in laboratory experiments that collisions can take place and induce significant wave energy dissipation. Smith and Thomson (2020) observed collisions between ice pancakes in the newly formed MIZ area. Rabault et al. (2019) used optical methods to observe collisions between ice chunks and the associated water pumping in a small scale laboratory experiment. This demonstrated that floe-floe collisions can cause wave-in-ice energy dissipation due to both the inelastic energy losses in the collisions themselves, as well as the increased water turbulence levels and effective eddy viscosity induced in the water immediately underneath the ice. Following these results, the study by Løken et al. (2022) found in idealized field experiments that floe-floe collisions can amount to up to 45 % of the energy dissipated when ice floes move close to each other, and confirmed that collisions, when occurring, are at the origin of complex water jets and eddies. Furthermore, floe-floe collisions are also believed to be strongly related to the floe size distribution and likely to play a role in the merging and breaking of ice floes (Shen and Ackley, 1991; Herman, 2018). These collisions are believed to attenuate incoming waves both through the dissipation of energy from the collisions themselves and through increased eddy viscosity resulting from the injected turbulent kinetic energy. The dissipation due to floe-floe collisions has also been modeled by Shen and Squire (1998), whose model found floe-floe collisions to be the predominant wave attenuation mechanism in compact pancake floe fields, and Rottier (1992), who developed one of the first models on floe collision rates. While the early observations by Martin and Becker (1987) and Martin and Becker (1988), are important, the MIZ and polar regions are undergoing rapid changes due to anthropogenic climate change (Kinda, 2013, chapter 5.1). The sea ice is getting thinner (going from a mean of 3.64 m to 1.89 m from 1980 to 2008) (Kwok and Rothrock, 2009), younger (Maslanik et al., 2007), and the MIZ constitutes an increasingly large part of the total polar ice sheets (Strong and Rigor, 2013). Hence, continued monitoring of the arctic ice sheets are important to adequately quantify if and how the dynamics in the area are changing with the climate. Yet, the amount of observations of the phenomenon of floe-floe collisions has been low in the recent years.Noyce et al. (2023) recently reported time series of floe-floe collisions in the outer Antarctic MIZ during a polar cyclone. Recently, Rabault et al. (2024) has presented observations that may indicate that sea ice convergence could modulate the intensity of floe-floe collisions, with large impact on wave in ice attenuation.
@@
Here, we provide data showing collisions between ice floes in the arctic MIZ, and discuss the causal mechanisms and circumstances surrounding these events. We find more than 60 acceleration residuals in the horizontal plane that lie outside the 99.9% probability region in two time series. This corresponds to more than five times the expected amount of extreme events, and is analogous to a $3\sigma$-event in the one dimensional setting. These collisions are aligned with the dominating wave direction, and phase locked with the wave signal, strongly suggesting that the individual collisions we detect are wave-induced.
@@ 2.1 Buoy deployment and raw data recovery
The OMB-v2018 performs wave measurements using a Vectornav VN100 Inertial Motion Unit (IMU), and uses GPS to determine its position. The wave motion data are sampled over 20-minute intervals every four hours from records of vertical acceleration sampled at 10 Hz by the VN100 (Rabault et al., 2017). The VN100 has built-in accelerometer, gyroscope, and magnetometer, and performs on-board Kalman filtering (Rabault et al., 2016, 2020). This filtering ensures that the full 3-dimensional acceleration experienced by the buoy in the North-East-Down (NED) frame of reference is captured in the time series. Due to the presence of large batteries and cables carrying electrical currents in close vicinity to the VN100, the magnetometer data come with significant uncertainty and orientation relative to north should only be considered as an indication. The accuracy of the VN100 IMU is $5\cdot 10^{-3}$ g, with g being the acceleration of gravity. A comprehensive list over the tests and validations done on the VN100 IMU on the OMB performed prior to deployment is given in section 2.1 in Rabault et al. (2020).
@@
Figure 2: Trajectories of the buoys deployed 2020-08 over the Yermak plateau area, superimposed on synthetic aperture radar (SAR) images of the MIZ on 2020-08-29. The buoy that was recovered is the northernmost one, which is shown in red. The markers indicate the position of each buoy on the 2020-08-29 at 12 UTC. The figure is created using the Ocean Visual Laboratory tool (permalink: https://odl.bzh/vEFWnkH0). Sea ice is shown in gray, with dark gray spots being larger ice floes, in the SAR images, and the buoy recovered is deep in the MIZ.
@@ 2.2 Signal processing of recovered in-situ time series
where $f_{\text{max}}$ is the frequency at which the PSD reaches its maximum value. Furthermore, we estimated the steepness of the wave field, we assumed for simplicity the dispersion relation to be that of deep-water waves (which holds true to leading order, see equations 11 and 12 in Meylan et al. (2018)). Hence, using the significant waveheight as the wave amplitude, and peak period as the wave period, we estimate the steepness to be:
$a_{k}=\frac{4\pi^{2}H_{s}}{gT_{p}^{2}}$ (5)
@@ 2.3 Detection of Collisions
The measured accelerations contain three contributions: (i) wave-induced motion of the ice floe, (ii) other physically meaningful signals, in particular, as we will demonstrate in the following, floe-floe collisions, (iii) sensor and signal processing (in particular Kalman filter) noise. The latter two, i.e. (ii) and (iii), correspond to the residual signal computed above.
The noise contribution (iii) arises from the noise in the micro-electromechanical systems (MEMS) sensors used in the IMU, and its propagation through the embedded Kalman filter run on-the-fly. According to the datasheet of the IMU, and in agreement with previous observations (Rabault et al., 2020), this generally follows a normal distribution. Hence, statistically significant deviations from a normal distribution in the residuals is a likely signature of acceleration events which can not be attributed to noise. This is tested using a Pearson $\chi^{2}$-test. Furthermore, we compute the 99% and 99.9% probability areas for the residuals under an assumption that the residuals are normally distributed. Using a visual inspection of the flagged files, we can check if an unexpectedly large amount of residuals fall outside these areas, and if so, the residuals are most likely not normally distributed and likely contain physically meaningful signals (ii).
@@ 2.4 Ocean and sea ice model data
Model data leveraging assimilation of satellite observations are used to estimate ocean and sea ice conditions in the area where the recovered buoy was active. In particular, we use the WAM-4km model data, run operationally at the Norwegian Meteorological Institute (MetNo), to estimate the open ocean wave conditions namely swell, wind speed and wind sea waves (see (Group, 1988; Günther et al., 1992; Ardhuin et al., 2010; Breivik et al., 2022)). The corresponding model data are openly available [^2] https://thredds.met.no/thredds/catalog/fou-hi/mywavewam4archive/catalog.html.
Estimates for sea ice concentration and sea ice thickness were gathered from the Nordic4 model ( see (Budgell, 2005; Spreen et al., 2008)), run operationally at MetNo and openly available [^3] https://thredds.met.no/thredds/catalog/fou-hi/nordic4km-zdepths1h/catalog.html.
@@ 3.1 Collisions detection: direction and phase locking with incoming waves
The extreme values in the residuals are likely to be caused by floe-floe collisions induced by the incoming waves. This is supported both by the general alignment of the residual acceleration extreme events in the horizontal plane with the dominating incoming wave direction, and the absence of extreme residuals in the vertical direction. In order to further cross-check this hypothesis, we look at the correlation between the instantaneous wave phase obtained by the Hilbert transform of the double integrated wave elevation $\eta(t)$, and the occurrence of extreme events in the acceleration residuals. If the extreme residual events are wave-induced collisions, we would expect the majority of the extreme events to happen at the the same point during the wave phase, as that point in the wave phase corresponds to the convergence of the colliding ice floes . As visible in Fig. 7, the extreme events are not uniformly distributed, with the majority of collisions in F291 happening in the third and fourth quadrants of the phase diagram and F312 in the second and third quadrants, and hence the extreme events follow the phase of the wave field. The extreme events of the control file, F335, are both not as extreme as the ones in the two FOIs, and do not exhibit any clear clustering in the data.
Figure 7: Phase diagram of the extreme residual values detected the two FOIs and the control file F335 where no collision like events were detected, relative to the incoming wave phase. The phase of the wave field is found using a Hilbert transform. Then, each extreme residual event (meaning it is outside the 99.9% probability region) is plotted at its respective phases. The upper half of the circle is the wave crest, while the bottom half is the wave trough. In the two FOI, we see signs of phase clustering, indicating that the occurrence of the extreme events we detect is strongly correlated to specific portions of the incoming wave phase motion. The extreme events in F335 are both less extreme (no $4\sigma$-events were observed) and less clustered in the phase diagram.
@@ 3.2 Waves and sea ice conditions when collisions happen
Figure 8: Sea ice concentration (SIC) encountered by waves propagating into the MIZ following the direction of the swell. The SIC is read from the ice model, while the closest available swell direction is read from the wave model. The blue dashed line shows the threshold corresponding to a SIC of 5%. The red dots indicate the buoy position along the swell direction relative to the edge of the MIZ. All SIC profiles are aligned to start right at the start of the MIZ (according to the 5% SIC threshold criterion) at the 0 km distance. The two FOIs and the control file F335 are highlighted in a dark color. The buoy is, in most files, more than 60 km from the edge of the MIZ alongside the propagating swell.
@@
Figure 9: Situation map showing the sea state for the two FOIs at the time when the collisions occurred. The sea ice concentration is shown in all four figures. The top row (figures (A),(B) and (C)) shows the model swell partition significant wave height and direction at the time of F291 (left), F312 (middle) and F335 (right), respectively, while the bottom row (figures (D), (E) and (F)) similarly shows the model wind sea partition significant wave height and direction for the two FOIs and control file. The position of the buoy at the time corresponding to each file is marked with a light blue star.
@@
Figure 10: Significant wave height (red), peak period (blue) and wave steepness $a_{k}$ (green) obtained from processing of the raw data files obtained from the SD card of the recovered buoy as a function of time. The black dashed lines indicate the times for which we detect collisions, while the gray dashed and dotted line is the control file F335, where no collisions were found.
@@ 4 Discussion
While we have convincing evidence that our records contain wave-induced floe-floe collisions, it is difficult to be more specific about the detailed mechanics of these collisions. While the motion of the sea ice under the influence of waves is mostly happening in the vertical direction, some residual horizontal motion can also take place. Therefore, if neighboring ice floes are close enough, it is possible that this residual horizontal motion can lead to floe-floe collisions, as recently discussed in Rabault et al. (2024). While one could expect that adjacent floes with similar shapes move in (near) synchronization under the influence of monochromatic waves, conditions in the field can lead to the existence of significant relative motion between floes. In particular, real world wave fields are not monochromatic and individual waves have a spread in both their individual amplitude and wavelength. This naturally leads to different forcing between consecutive waves, which can result in different motion between adjacent floes at any given time, and, hence, collisions. In addition, ice floe sizes typically follow a relatively broad floe size distribution (Toyota et al., 2006). This means that adjacent floes can have different responses to the same incoming waves, which, in turn, can also be a source of collisions. However, more work will be needed to investigate the detailed underlying mechanics, possibly by combining buoy observations with other more exhaustive techniques.
@@
The absence of more buoy measurements also means we are not able discuss the attenuation of the wave field in details. Furthermore, we are not able to discuss the prevalence of collisions in details either, as while we have found evidence of floe-floe collisions in two files, this is not a sufficiently large and well-sampled dataset to extrapolate our results. In the best case scenario, the data might give an estimate of the collision rate of an ice floe of the same size (which is unknown) as the floe at which the buoy was deployed. But more critically, the data we have indicate the collisions to be wave-induced. Hence, when over half the measurements were taken during conditions with little to no wave activity (see figure 10), any discussion on the frequency of collisions and their possible attenuating effect is challenging from these data alone. However, we want to stress that our present observations fit well with previous results. Løken et al. (2022) ran experiments where two ice floes were rammed into each other using winches, and studied the generation of TKE from the collisions and found that $36.9\%\pm 23.7\%$ of the input energy was dissipated through turbulence. Furthermore Herman et al. (2019) found the wave attenuation induced by colliding ice floes to be proportional to the square of the orbital velocity underneath. A model for the attenuation of ice floes was published by Shen and Squire (1998), in which a theoretical model for the attenuation by collisions between pancake floes is derived. Our results can not be directly compared to their model, as we are not able to infer the attenuation rate of the wave field and the floe size and distribution in the area. Yet, we still observe that Shen and Squire (1998) found a clear wave amplitude dependency on the collision frequency, with all three model runs predicting low collision frequencies $f<<1$ Hz when the wave amplitude is low enough. We do, however, still note that the model of Shen and Squire (1998) assumes pancake floes at a given size.
The present study owes a lot to luck and serendipity. Both the fact that the instrument v-2018 contains a SD card with a full copy of the time series (which is due to the fact that the microcontrollers available back then were not able to perform the signal processing needed, so that an additional RaspberryPi had to be added for processing the time series), and the fact that the instrument recovered got stranded, was found, and that the SD card was not damaged, are a strike of luck. However, it would be easily possible, based on the experience obtained from the present dataset, to adapt existing open source instruments firmware, such as the OMB-v2021, to i) save raw time series on an embedded SD card in case the instrument is found again, ii) compute in-situ a number of collision statistics, such as the ones used here, and append these to the wave spectra that are being transmitted. Since a simple metrics of collisions, such as the number and relative strength of extreme acceleration events and their correlation with the wave phase, or a deviation in the distribution from the expected distribution of wave-induced acceleration measurements (as implemented by e.g. Brown et al. (2017)), would be only a few scalars, this would not add prohibitive transmission costs nor energy consumption to the design. Moreover, all the data needed to perform such an analysis are already available on-board. Adding such metrics to wave in ice buoys transmissions would allow to perform large scale, statistically representative analysis of the correlation between collision-like proxy metrics, changes in the sea ice and wave in ice conditions, and the wave in ice attenuation rate.